OpenStudy (anonymous):

whats the LCD 5x/7+1/14=7/28?

jhonyy9 (jhonyy9):

so there are denominators 7 =7*1 14 =7*2 28 =7*4

jhonyy9 (jhonyy9):

so than now if you check these how do you think what will be the low common denominator ?

jhonyy9 (jhonyy9):

do you know what mean LCD ?

jhonyy9 (jhonyy9):

ok so i think the LCD there will be 28 do you know why will be the 28 right ?

jhonyy9 (jhonyy9):

so because there the denominator of the first fraction is 7 and you can getting the 28 if you multiplie the 7 by 4 - the denominator of the second fraction is there 14 and if you check it will get that you can getting easy the 28 if you multiplie the 14 by 2 - for the 3rd fraction is the denominator 28 ,so here i think not mean problem nothing ,because you know that if you multiplie the 28 by 1 you will get 28 so from these all you have got that for denominators 7,14 and 28 the LCD will be 28 sure hope so much that you understand these all yw good luck bye

OpenStudy (anonymous):

thanks